OpenStudy (anonymous):

Okay since I don't really know how to use this I'll just post the equation in the reply box. Express y in x.

OpenStudy (anonymous):

\[\log_{2} xy = 2 + 3 \log_{2} y - \log_{2} y\]

OpenStudy (anonymous):

(all the logs have a base of 2) logxy=2+3logy-logy logx+logy=log(4)+log(y^3)-logy logx-log4=log(y^3)-2logy log(x/4)=log(y^3)-2log(y^2) log(x/4)=log((y^3)/(y^2)) log(x/4)=log(y) y=x/4